This is a factoring problem with that may of may not have more than one solution. I need to see the steps to resolving it and any/all solutions. Can anyone help me?
w³ + 5w² - w = 5


Thanks

Answer by Nate(3500)

Use Descarte's Rule of Signs and the general works to get the answer:
w³ + 5w² - w - 5 = 0
(x - 1)(x^2 + 6x + 5) = 0
(x - 1)(x + 5)(x + 1) = 0
x = 1
x = -5
x = -1